How do I multiply 7(r+2)(r-2)
Answer by checkley71(8403)

7(r+2)(r-2) or using the FOIL (multiply the first 2 terms r*r=r~2, multiply the outer terms -2*r=-2r, multiply the inner terms +2*r=2r then multiply the last two terms 2*-2=-4 thus we have (r~2-2r+2r-4) or r~2-4 now we multiply this by 7(r~2-4)=7r~2-28